the perimeter of a rectangle is 280 cm. the ratio of the width to the length is 3:4. what is the length of the rectangle?

Answer by stanbon(75887)

the perimeter of a rectangle is 280 cm. the ratio of the width to the length is 3:4. what is the length of the rectangle?
-----
Perimeter = 2(length + width)
----
3:4 is the same as 3x:4x where 3x is width and 4x is length.
----
Equation:
280 = 2(4x+3x)
140 = 7x
x = 20
----
width = 3x = 3*20 = 60 cm
length = 4x = 4*20 = 80 cm
---------------------------
Cheers,
Stan H.
----------